Vehicles often travel city streets adjacent to parking lanes at 56 kph or faster. At his speed and setting detection through response-initiation time for an alert driver at 2 sec. and f = 0.50, how far must the driver be away from a suddenly opened car door to avoid striking it?


1
Expert's answer
2021-10-11T07:01:21-0400

"SSD = V_t+\frac{v^2}{2gf}\\
SSD =56*\frac{2}{18}+\frac{(56*\frac{2}{18})^2}{2*9.81*0.5}\\
SSD=10.16881 m"